OpenStudy (anonymous):

need to find the equation of the line containing the given pair of points (-2,-1) (-6,-7)

OpenStudy (anonymous):

ok this is how you do it

OpenStudy (anonymous):

you know the slope formula?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

that the regular formula

OpenStudy (anonymous):

m=(y-y1)/(x-x1) is the slope

OpenStudy (anonymous):

o that one

OpenStudy (anonymous):

plug in your value tell me what you get?

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (anonymous):

then you mimic the slope formula

OpenStudy (anonymous):

m=(-1-(-7)/-6-(-2)

OpenStudy (anonymous):

yes, so what you get?

OpenStudy (anonymous):

6/4 is your slope, positive slope

OpenStudy (anonymous):

you know what to do next?

OpenStudy (anonymous):

um im drawing a blank

OpenStudy (anonymous):

y-y1 = m(x-x1) you bring the x over then +y1 to the right side and you get y=y1+m(x-x1)

OpenStudy (anonymous):

or rather its just F(x)=y1+m(x-x1)

OpenStudy (anonymous):

you get your slope? 6/-4

OpenStudy (anonymous):

sorry its was negative slope

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

F(x)=y1+m(x-x1) is the equation of a line

OpenStudy (anonymous):

F(x)=y1+(6/-4)(x-x1)

OpenStudy (anonymous):

you plug in either or any of the points for x1,y1 solve for the equation and there you go eq of a line

OpenStudy (anonymous):

both will get you the same answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

eq of a line is very simple just remember the slope formula bring the denominator over to the m and isolate y itself

OpenStudy (anonymous):

math is my weak subject

OpenStudy (anonymous):

dont worry its mine also i am trying to solidify my understanding in math, i am having trouble in calculus

OpenStudy (anonymous):

i can do everything but i cant do math

OpenStudy (anonymous):

hahaha same, lol i was just refreshing up my algebra, i came across this site last few days

OpenStudy (anonymous):

but take your time reading my steps i put up it should be clear

OpenStudy (anonymous):

yeah me too!! Its a lifesaver. I have another one too

OpenStudy (anonymous):

cool you can add me as fan or if you have other questions i'll be glad to help i'm like doing math at 2 am in the morning lol in the summer

OpenStudy (anonymous):

yeah ive been doing the same thing lately. Im cramming because im leaving for Vegas on Monday and need to get this done

OpenStudy (anonymous):

r u in the US? If so what time zone?

OpenStudy (anonymous):

eastern

OpenStudy (anonymous):

ok, Im Pacific

OpenStudy (anonymous):

take care i am dead tired i gonna sleep now

OpenStudy (anonymous):

thanks for all ur help and im sure u will hear from me again. Sleep well